# Distributionally Robust Optimization with Adversarial Data Contamination

### v1

@@ 2 Preliminaries
Let $l:{\mathbb{R}}\to{\mathbb{R}}$ be a convex function. The convex conjugate (or Fenchel conjugate) of $g$ is the function $l^{*}:{\mathbb{R}}\to{\mathbb{R}}\cup\{+\infty\}$ defined by $l^{*}(\alpha)\coloneqq\sup_{\hat{y}\in{\mathbb{R}}}\{\alpha\hat{y}-l(\hat{y})\}$.
@@
Robust mean estimation
Recent developments in robust mean estimation algorithms achieve dimension-independent error guarantees under the strong contamination model. A family of such algorithms rely on the notion of $(\epsilon,\delta)$-sample stability [diakonikolas2022algorithmic, Definition 2.1].
@@
The following characterization of stability condition is convenient when (uncorrupted) data generating distribution ${\mathbb{P}}_{0}$ is only assumed to have bounded operator norm of its covariance matrix:
Fact 2.6 (Stability under bounded covariance [diakonikolas2022algorithmic, Lemma 3.11]).
@@
One such algorithm is provided in Appendix B. State-of-the-art algorithms for robust mean estimation can be implemented in near-linear time, requiring only a logarithmic number of passes over the data; see, e.g., [diakonikolas22streaming, ChengDG19, dong2019quantum]. Any such faster algorithm could be adopted for our purposes.
Fact 2.7 (Robust Mean Estimation with Stability [diakonikolas2020outlier, Theorem A.3]).
@@
Fact 2.7 requires uncorrupted samples to be stable. With a sufficiently large sample size, most samples from a bounded covariance distribution are stable. The remaining samples can be treated as outliers without an effect on the (order of the) estimation error.
Fact 2.8 (Sample Complexity for Stability [diakonikolas2022algorithmic, Proposition 3.9] and [diakonikolas2020outlier, Theorem 1.4]).
Fix $\tau\in(0,1)$. Let $S$ be a set of $N$ independent samples from a distribution on $\mathbb{R}^{d}$ with mean ${\bm{\mu}}$ and covariance ${\bm{\Sigma}}$. With probability at least $1-\tau$, the subset $S^{\prime}=\{{\bm{x}}\in S:\|{\bm{x}}-{\bm{\mu}}_{S}\|_{2}\leq 2\sqrt{\|{\bm{\Sigma}}\|_{\operatorname{op}}}\sqrt{d/\epsilon}\}$ satisfies $\left|S^{\prime}\right|\geq\left(1-\epsilon\right)|S|$ and $S^{\prime}$ is $\left(\epsilon,O(\sqrt{\epsilon})\right)$-stable with respect to ${\bm{\mu}}$ and $\|{\bm{\Sigma}}\|_{\operatorname{op}}$, provided the sample size $N=O\big{(}(d\log d+\log(1/\tau))/{\epsilon}\big{)}$ is sufficiently large.
@@ 3 Data-contaminated DRO: Modeling Principles and Problem Formulation
We begin this section with an overview of our main modeling principles and a review of most closely related work, pointing out important differences. We then argue that under standard assumptions about the loss function and the reference (generating) distribution, we can reduce the considered problem to a convex optimization problem with an inexact data oracle.
@@ 3.1 Modeling Principles
The modeling principles follow similar considerations as [blanchet2024distributionally]. Let ${\mathbb{P}}_{0}$ denote the data generating distribution (also called the reference distribution): this is the distribution from which clean, uncorrupted training samples would be drawn if there were no data corruptions. The training samples passed to a learning algorithm might contain outliers, be contaminated, or otherwise have discrepancy with ${\mathbb{P}}_{0}$; thus we let ${{\mathbb{P}}_{\text{train}}}$ denote the contaminated distribution that generates the training data. A sample set of size $N$ is then drawn from ${{\mathbb{P}}_{\text{train}}}$, leading to the empirical distribution ^=^${\widehat{\mathcal{{\mathbb{P}}}}}={\widehat{\mathcal{{\mathbb{P}}}}}(N)$. The learner has access to ^${\widehat{\mathcal{{\mathbb{P}}}}}(N)$, based on which it outputs a model parameterized by a vector ${\bm{w}}$. The model parameterized by ${\bm{w}}$ is tested against an out-of-sample distribution (also known as testing distribution) ${{\mathbb{P}}_{\text{test}}}$, which may deviate from ${\mathbb{P}}_{0}$. This process can be summarized as:
[…]
The model parameterized by ${\bm{w}}$ might perform poorly on ${{\mathbb{P}}_{\text{test}}}$ for various reasons: (i) Overfitting/statistical error: ^${\widehat{\mathcal{{\mathbb{P}}}}}(N)\neq{{\mathbb{P}}_{\text{test}}}$; (ii) Distributional shift: ${{\mathbb{P}}_{\text{test}}}\neq{\mathbb{P}}_{0}$; or (iii) Data contamination: ${\mathbb{P}}_{0}\neq{{\mathbb{P}}_{\text{train}}}$. Errors due to (i) and (ii) arise in the post-decision stage, while the error in (iii) is incurred in the pre-decision stage. Distributionally robust optimization (DRO) minimizes the loss with respect to distributions in some uncertainty (or “ambiguity”) set that is specified without inspecting the training data; as a consequence, it is best used to model post-decision errors. In this work, we specifically focus on DRO’s advantage in guarding against distributional shifts. We gloss over the overfitting errors by drawing samples with a sufficiently large sample set size $N$ and apply uniform convergence analysis to guarantee the performance of every model ${\bm{w}}$ on ${{\mathbb{P}}_{\text{train}}}$.
@@ 3.2 Related Models
In prior work [nietert2022outlier, nietert2023outlier], outlier robustness in addition to Wasserstein distributional robustness was handled by introducing a new distance measure
@@
We remark that (OR-WDRO) seems to be able to also model errors from (iii) data contamination, as claimed in [nietert2023outlier], in the following sense: suppose the training data are corrupted so that $\operatorname{TV}({{\mathbb{P}}_{\text{train}}},{\mathbb{P}}_{0})\leq\epsilon$; then ${\mathbb{P}}_{0}\in\{{\mathbb{Q}}:W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{Q}})=0\}$. If, in addition, we allow distribution shifts and assume that $W_{p}({{\mathbb{P}}_{\text{test}}},{\mathbb{P}}_{0})\leq\rho$, then $W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{{\mathbb{P}}_{\text{test}}})\leq\rho$. In this sense, (OR-WDRO) minimizes the worst-case loss with respect to an ambiguity set that contains ${{\mathbb{P}}_{\text{test}}}$. However, we argue that modeling data contamination with (OR-WDRO) has multiple drawbacks; see Section C.1 for details.
@@
    Dimension dependence: the excess risk bound of the solution of (OR-WDRO) depends on the dimension of the data ${\bm{\xi}}$ (or the dimension of the underlying features), which can be excessive in high-dimensional settings [nietert2023outlier, Theorems 1 and 4].
@@
    We illustrate that (OR-WDRO) cannot be used to formulate the most basic task in robust statistics—robust mean estimation—under Wasserstein-$p$ distributional shifts with $p<2$.
@@
    We prove that the limit behavior of (OR-WDRO) as $\rho\to 0$ is a variant of Conditional Value at Risk (CVaR) [^3] The limit behavior of the tractable formulation in [nietert2023outlier, Section 3.3] is exactly CVaR under limits; see Lemma C.3., which is known to be sensitive to outliers; see Theorem 2.1 and the discussion thereafter in [blanchet2024distributionally], noting that CVaR can be interpreted as the limit of some $f$-divergence DRO risks [duchi2021learning, Example 3]. In other words, a model parameterized by the optimal solution to (OR-WDRO) is not robust to outliers in the training set, even without any distributional shifts [^4] In fact, the maximizing distribution $\operatorname*{arg\,max}_{{\mathbb{Q}}}\big{\{}\mathbb{E}_{{\bm{\xi}}\sim{\mathbb{Q}}}[\ell({\bm{w}};{\bm{\xi}})]:\exists{\mathbb{P}}^{\prime},\epsilon^{\prime}\leq\epsilon\text{ s.t. }{{\mathbb{P}}_{\text{train}}}=(1-\epsilon^{\prime}){\mathbb{Q}}+\epsilon^{\prime}{\mathbb{P}}^{\prime}\big{\}}$ puts weights only on $(1-{\epsilon})$-fraction of samples with largest loss values, so in this sense CVaR-DRO is the opposite of robust estimation with data contamination..
@@
${{\mathbb{P}}_{\text{train}}}={\mathbb{P}}_{0}$ $W_{p,\epsilon}({{\mathbb{P}}_{\text{test}}},{\mathbb{P}}_{0})\leq\rho$ OR-WDRO [nietert2023outlier]
$W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{P}}_{0})\leq\rho$ [^5] Training data in [pittas2024optimal] can simultaneously be ${\epsilon}$-corrupted and have Wasserstein perturbation. ${{\mathbb{P}}_{\text{test}}}={\mathbb{P}}_{0}$ Local & global corruption [nietert2024robust, pittas2024optimal]
@@
Another approach that, like ours, simultaneously addresses both training data contamination (modeled via $\operatorname{TV}({\mathbb{P}}_{0},{{\mathbb{P}}_{\text{train}}})\leq{\epsilon}$) and distributional shifts at test time is the Distributionally Robust Outlier-aware Optimization (DORO) framework proposed by [zhai2021doro]. DORO defines its objective as the minimum $f$-divergence-based DRO risk (with shift parameter $D_{f}({{\mathbb{P}}_{\text{test}}},{{\mathbb{P}}_{\text{train}}})\leq\rho$) among all possible $(1-\epsilon)$-subpopulations of the training data. While the authors provide theoretical guarantees for the DORO risk minimizer, their proposed algorithm lacks both error bounds and computational complexity guarantees; as such, it remains unclear whether the DORO risk minimizer can be approximated provably—let alone efficiently (see further discussion in Section C.2). Furthermore, the theoretical guarantees of the DORO risk are predicated on unconventional assumptions and thus do not recover classical outlier-robust learning guarantees when the distributional shift parameter $\rho$ is set to zero.
The work by [wang2024outlier] introduces a DRO framework defined by a novel ambiguity measure inspired by Unbalanced Optimal Transport (UOT). They propose that UOT-based distance is inherently more resilient to outliers and their formulation subtracts a penalty term based on prior knowledge to further down-weight contaminated data. While [wang2024outlier] used this approach to handle contaminated training data and provided empirical results demonstrating improved robustness in such settings compared to standard Wasserstein DRO and OR-WDRO, they did not establish explicit theoretical guarantees on estimation error. More importantly, from the modeling perspective of (2), its mechanism of defining an ambiguity suggests that the outliers it robustifies against are perhaps best characterized in the testing distribution (${{\mathbb{P}}_{\text{test}}}$), as discussed previously for [nietert2023outlier] and in Section 3.1.
In the remaining text, we use the word data contamination or corruptions to specifically indicate the existence of outliers in the training data, whereas outliers might exist in either training or testing data.

### v2

@@ 2.1 Optimization
[…]
@@
2.2 Robust mean estimation
Recent developments in robust mean estimation algorithms achieve dimension-independent error guarantees under the strong contamination model. A family of such algorithms rely on the notion of $(\epsilon,\delta)$-sample stability [DK23a, Definition 2.1].
@@ 2.2 Robust mean estimation
The following characterization of the stability condition is convenient when (uncorrupted) data generating distribution ${\mathbb{P}}_{0}$ is only assumed to have bounded operator norm of its covariance matrix:
Fact 2.6 (Stability under bounded covariance [DK23a, Lemma 3.11]).
@@
One such algorithm is provided in Appendix˜A. State-of-the-art algorithms for robust mean estimation can be implemented in near-linear time, requiring only a logarithmic number of passes over the data; see, e.g., [Dia+22a, CDG19a, DHL19a]. Any such faster algorithm could be adopted for our purposes.
Fact 2.7 (Robust Mean Estimation with Stability [DKP20a, Theorem A.3]).
@@
Fact˜2.7 requires uncorrupted samples to be stable. With a sufficiently large sample size, most samples from a bounded covariance distribution are stable. The remaining samples can be treated as outliers without an effect on the (order of the) estimation error.
Fact 2.8 (Sample Complexity for Stability [DK23a, Proposition 3.9] ).
Let $S$ be a set of $N$ independent samples from a distribution on $\mathbb{R}^{d}$ with mean $\bm{\mu}$} and covariance $\bm{\Sigma}$}. Fix $\tau\in(0,1)$. With probability at least $1-\tau$, the subset $S^{\prime}=\{{\bm{x}}\in S:\|{\bm{x}}-{\bm{\mu}}_{S}\|_{2}\leq 2\sqrt{\|{\bm{\Sigma}}\|_{\operatorname{op}}}\sqrt{d/\epsilon}\}$ satisfies $\left|S^{\prime}\right|\geq\left(1-\epsilon\right)|S|$ and $S^{\prime}$ is $\left(\epsilon,O(\sqrt{\epsilon})\right)$-stable with respect to $\bm{\mu}$} and $\|{\bm{\Sigma}}\|_{\operatorname{op}}$, provided the sample size $N=O\big((d\log d+\log(1/\tau))/{\epsilon}\big)$ is sufficiently large.
@@ 3 Data-contaminated DRO: Modeling Principles and Problem Formulation
In this section, we develop our conceptual framework for addressing data contamination and distributional shifts simultaneously. We begin by establishing our main modeling principles, followed by a critical review of the most closely related models to motivate our specific formulation. We then argue that under standard assumptions about the loss function and the reference (generating) distribution, we can reduce the considered problem to a convex optimization problem with an inexact data oracle.
@@ 3.1 Modeling Principles
The modeling principles follow similar considerations as [Bla+25a]. Let ${\mathbb{P}}_{0}$ denote the data generating distribution (also called the reference distribution): this is the distribution from which clean, uncorrupted training samples would be drawn if there were no data corruptions. The training samples passed to a learning algorithm might contain outliers, be contaminated, or otherwise have discrepancy with ${\mathbb{P}}_{0}$; thus we let ${\mathbb{P}}_{\text{train}}$} denote the contaminated distribution that generates the training data. A sample set of size $N$ is then drawn from ${\mathbb{P}}_{\text{train}}$}, leading to the empirical distribution ${\widehat{\mathcal{{\mathbb{P}}}}}={\widehat{\mathcal{{\mathbb{P}}}}}(N)$. The learner has access to ${\widehat{\mathcal{{\mathbb{P}}}}}(N)$, based on which it outputs a model parameterized by a vector $\bm{w}$}. The model parameterized by $\bm{w}$} is tested against an out-of-sample distribution (also known as testing distribution) ${\mathbb{P}}_{\text{test}}$}, which may deviate from ${\mathbb{P}}_{0}$. This process can be summarized as:
${\mathbb{P}}_{0}\xrightarrow[]{\text{contamination}}{{\mathbb{P}}_{\text{train}}}\xrightarrow[]{\text{sampling}}{\widehat{\mathcal{{\mathbb{P}}}}}(N)\xrightarrow[]{\text{train-to-test shift}}{{\mathbb{P}}_{\text{test}}}.$ (2)
The model parameterized by $\bm{w}$} might perform poorly on ${\mathbb{P}}_{\text{test}}$} for various reasons: (i) Overfitting/statistical error: ${\widehat{\mathcal{{\mathbb{P}}}}}(N)\neq{{\mathbb{P}}_{\text{test}}}$; (ii) Distributional shift: ${{\mathbb{P}}_{\text{test}}}\neq{\mathbb{P}}_{0}$; or (iii) Data contamination: ${\mathbb{P}}_{0}\neq{{\mathbb{P}}_{\text{train}}}$. Errors due to (i) and (ii) arise in the post-decision stage, while the error in (iii) is incurred in the pre-decision stage. Distributionally robust optimization (DRO) minimizes the loss with respect to distributions in some uncertainty (or “ambiguity”) set that is specified without inspecting the training data; as a consequence, it is best used to model post-decision errors. In this work, we specifically focus on DRO’s advantage in guarding against distributional shifts. We gloss over the overfitting errors by drawing samples with a sufficiently large sample set size $N$ and apply uniform convergence analysis to guarantee the performance of every model $\bm{w}$} on ${\mathbb{P}}_{\text{train}}$}.
@@ 3.2 Related Models
In prior work [NGC22a, NGS23a], outlier robustness in addition to Wasserstein distributional robustness was handled by introducing a new distance measure
@@
We remark that (OR-WDRO) seems to be able to also model errors from training data contamination (i.e., Case (iii)) in the following sense: suppose the training data are corrupted so that $\operatorname{TV}({{\mathbb{P}}_{\text{train}}},{\mathbb{P}}_{0})\leq\epsilon$; then ${\mathbb{P}}_{0}\in\{{\mathbb{Q}}:W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{Q}})=0\}$. If, in addition, we allow distribution shifts and assume that $W_{p}({{\mathbb{P}}_{\text{test}}},{\mathbb{P}}_{0})\leq\rho$, then $W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{{\mathbb{P}}_{\text{test}}})\leq\rho$. In this sense, (OR-WDRO) minimizes the worst-case loss with respect to an ambiguity set that contains ${\mathbb{P}}_{\text{test}}$}. However, we argue that modeling data contamination with (OR-WDRO) has multiple drawbacks:
@@
    Dimension dependence: the excess risk bound of the solution of (OR-WDRO) depends on the dimension of the data $\bm{\xi}$} (or the dimension of the underlying features), which can be a significant factor in high-dimensional settings [NGS23a, Theorems 1 and 4].
@@
    For the canonical task of robust mean estimation, the (OR-WDRO) objective becomes ill-posed (evaluates to infinity) for Wasserstein-$p$ metrics with $p<2$.
@@
    We prove that the limit behavior of (OR-WDRO) as $\rho\to 0$ is a variant of Conditional Value at Risk (CVaR) [^4] The limit behavior of the tractable formulation in [NGS23a, Section 3.3 ] is exactly CVaR under limits; see Lemma 3.4., which is known to be sensitive to outliers; see Theorem 2.1 and the discussion thereafter in [Bla+25a], noting that CVaR can be interpreted as the limit of some $f$-divergence DRO risks [DN21a, Example 3]. In other words, a model parameterized by the optimal solution to (OR-WDRO) is not robust to outliers in the training set, even without any distributional shifts [^5] In fact, the maximizing distribution $\operatorname*{arg\,max}_{{\mathbb{Q}}}\big\{\mathbb{E}_{{\bm{\xi}}\sim{\mathbb{Q}}}[\ell({\bm{w}};{\bm{\xi}})]:\exists{\mathbb{P}}^{\prime},\epsilon^{\prime}\leq\epsilon\text{ s.t. }{{\mathbb{P}}_{\text{train}}}=(1-\epsilon^{\prime}){\mathbb{Q}}+\epsilon^{\prime}{\mathbb{P}}^{\prime}\big\}$ puts weights only on $(1-{\epsilon})$-fraction of samples with largest loss values, so in this sense CVaR-DRO is the opposite of robust estimation with data contamination..
The first implication follows directly from the cited work. We now elaborate on the latter two points with detailed technical arguments. We begin with a standard fact (folklore; proof provided in Section˜B.2) that we will invoke repeatedly.
Fact 3.1.
$\lim_{|z|\to\infty}\frac{\ell_{y}(z)}{|z|^{p+\iota}}=C$
for some constants $\iota>0$ and $C>0$ that are independent of $y$.
Then, $\sup_{{\mathbb{Q}}:W_{p}({\mathbb{P}}_{0},{\mathbb{Q}})\leq\rho}\mathbb{E}_{({\bm{X}}^{\prime},Y^{\prime})\sim{\mathbb{Q}}}[\ell_{Y^{\prime}}({\bm{w}}\cdot{\bm{X}}^{\prime})]=+\infty,$ where the Wasserstein distance $W_{p}$ is defined with respect to the cost function $c(({\bm{x}}_{1},y_{1}),({\bm{x}}_{2},y_{2}))=\|{\bm{x}}_{1}-{\bm{x}}_{2}\|+\kappa|y_{1}-y_{2}|$ for some $\kappa\in[0,+\infty]$.
We illustrate the second drawback through a canonical problem in robust statistics—robust mean estimation. In this setting, applying the (OR-WDRO) formulation in an attempt to directly account for outliers or corrupted samples in the training data often leads, under common conditions, to objectives that are uninformative or ill-posed.
Example 3.2 (Mean Estimation and OR-WDRO for Training Data Contamination).
Suppose we have training data from ${\mathbb{P}}_{\text{train}}$}, which contains both “local” Wasserstein-$1$ perturbation and “global” TV contamination, i.e., there exists a distribution $\mathbb{P}$$^{\prime}$ with $W_{p}({\mathbb{P}}^{\prime},{\mathbb{P}}_{0})\leq\rho$ and $\operatorname{TV}({\mathbb{P}}^{\prime},{{\mathbb{P}}_{\text{train}}})\leq\epsilon$. To use (OR-WDRO) formulation to define an ambiguity set of radius $\rho$ around ${\mathbb{P}}_{\text{train}}$} that is intended to contain ${\mathbb{P}}_{0}$, we set $W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{P}}_{0})\leq\rho$. Our goal in this example is to illustrate that such a modeling approach for training data contamination using OR-WDRO can be problematic for $p<2$. Therefore, for this example, we set the final test distribution of interest to be ${{\mathbb{P}}_{\text{test}}}={\mathbb{P}}_{0}$, as our aim is to scrutinize how OR-WDRO might model training data contamination that separates an observed corrupted distribution ${\mathbb{P}}_{\text{train}}$} from ${\mathbb{P}}_{0}$. The OR-WDRO objective is
which equals infinity when $p<2$ as a consequence of Fact˜3.1; see the proof and intuition below.
This highlights a modeling limitation of OR-WDRO when interpreted as a direct mechanism for handling training data contamination, because methods such as those discussed in [NGS24a] and [PP24a, Theorem 1.4 ] can provide finite error guarantees ($O(\sigma\sqrt{\epsilon}+\rho)$) for mean estimation when dealing with $\epsilon$-corrupted samples from some $\tilde{{\mathbb{P}}}$ such that $W_{1}(\tilde{{\mathbb{P}}},{\mathbb{P}}_{0})\leq\rho$.
We now provide proof that Equation˜3 equals $\infty$ when $p<2$.
Proof.
The ambiguity set ${\mathcal{U}}=\{{\mathbb{Q}}:W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{Q}})\leq\rho\}$ can be expanded using the definition of $W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{Q}})=\inf_{\operatorname{TV}({\mathbb{P}}^{\prime},{{\mathbb{P}}_{\text{train}}})\leq\epsilon}W_{p}({\mathbb{P}}^{\prime},{\mathbb{Q}})$. Thus, ${\mathcal{U}}=\{{\mathbb{Q}}:\exists{\mathbb{P}}^{\prime}\text{ s.t. }\operatorname{TV}({\mathbb{P}}^{\prime},{{\mathbb{P}}_{\text{train}}})\leq\epsilon\text{ and }W_{p}({\mathbb{P}}^{\prime},{\mathbb{Q}})\leq\rho\}$. This means that the ambiguity set $\mathcal{U}$} is the union of all $W_{p}$-balls of radius $\rho$ centered at every distribution $\mathbb{P}$$^{\prime}$ within the $\epsilon$-TV ball around ${\mathbb{P}}_{\text{train}}$}, i.e., ${\mathcal{U}}=\bigcup_{{\mathbb{P}}^{\prime}:\operatorname{TV}({\mathbb{P}}^{\prime},{{\mathbb{P}}_{\text{train}}})\leq\epsilon}B_{W_{p}}({\mathbb{P}}^{\prime},\rho)$. The OR-WDRO objective can then be expressed as:
Now consider any distribution $\mathbb{P}$$^{\prime}$ such that $\operatorname{TV}({\mathbb{P}}^{\prime},{{\mathbb{P}}_{\text{train}}})\leq\epsilon$. We know that $\mathbb{P}$$^{\prime}$ can have unbounded support because $\operatorname{TV}$ contamination is blind to geometry. Thus the inner supremum, $\sup_{{\mathbb{Q}}:W_{p}({\mathbb{P}}^{\prime},{\mathbb{Q}})\leq\rho}\mathbb{E}_{{\bm{\xi}}\sim{\mathbb{Q}}}[\|{\bm{w}}-{\bm{\xi}}\|_{2}^{2}]$, will be infinite for any ${\bm{w}}\in\mathbb{R}^{d}$ when $p<2$. This is a standard result (Fact˜3.1) for $W_{p}$-DRO with $p<2$ when the loss function (here, quadratic) grows faster than the $p$-th power of the transport cost; adversarial perturbations $\mathbb{Q}$} of $\mathbb{P}$$^{\prime}$ can shift mass to points $\bm{x}$} with arbitrarily large $\|{\bm{x}}\|_{2}$ at a bounded $W_{p}$ cost, driving the expected loss to infinity. ∎
We next analyze the limiting behavior of the OR-WDRO objective when the Wasserstein radius $\rho$ approaches zero. This analysis highlights an inherent connection to Conditional Value-at-Risk (CVaR)-like measures. Given that CVaR is known to be sensitive to extreme data points [^6] See Theorem 2.1 and the discussion thereafter in [Bla+25a]; CVaR can be interpreted as the limit of some $f$-divergence DRO risks [DN21a, Example 3]., this connection underscores why the OR-WDRO model may not achieve robust protection against training data outliers in the way dedicated robust statistics methods do.
Lemma 3.3 (Limit Behavior of OR-WDRO Formulation).
For any ${\bm{w}}\in\mathbb{R}^{d}$,
Proof.
By [BV22a, Theorem 2.4 ], it holds that
where $\operatorname{LP}_{\{{\bm{0}}\}}({\mathbb{P}},{\mathbb{Q}})$ is defined to be $\inf_{\pi\in\Pi({\mathbb{P}},{\mathbb{P}}^{\prime})}\int\mathbb{I}_{{\bm{\xi}}-{\bm{\xi}}^{\prime}\not\in\{{\bm{0}}\}}\mathrm{d}\pi(\xi,\xi^{\prime})$ and $\Pi({\mathbb{P}},{\mathbb{P}}^{\prime})$ denotes the set of joint distributions whose marginal distributions are $\mathbb{P}$} and $\mathbb{P}$$^{\prime}$; [Str65a] shows that $\operatorname{LP}_{\{{\bm{0}}\}}({\mathbb{P}},{\mathbb{P}}^{\prime})=\operatorname{TV}({\mathbb{P}},{\mathbb{P}}^{\prime})$. ∎
For tractable computation, [NGS23a] introduce a tractable variant of (OR-WDRO) using a one-sided discrepancy $W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}}\|{\mathbb{Q}})=\inf_{{\mathbb{P}}^{\prime}:{\mathbb{P}}^{\prime}\leq\frac{1}{1-\epsilon}{{\mathbb{P}}_{\text{train}}}}W_{p}({\mathbb{P}}^{\prime},{\mathbb{Q}})$ (where $\mathbb{P}$$^{\prime}$ is a probability measure) and imposing a moment constraint ${\mathbb{Q}}\in{\mathcal{G}}_{2}(\sigma,{\bm{z}}_{0}):=\{\pi:\mathbb{E}_{{\bm{\xi}}\sim\pi}[\|{\bm{\xi}}-{\bm{\xi}}_{0}\|_{2}^{2}]\leq\sigma^{2}\}$. Their tractable objective is:
Lemma 3.4 (Limit Behavior of Tractable OR-WDRO Reformulation).
Proof.
Based on the dual formulation in [NGS23a, Remark 3 ]:
As $\sigma\to+\infty$, the term $\lambda_{1}\sigma^{2}$ forces the optimal $\lambda_{1}$ to be 0 to maintain a finite infimum. Thus, the expression converges to $\operatorname{CVaR}_{{\bm{\xi}}\sim{{\mathbb{P}}_{\text{train}}}}^{1-\epsilon}[\ell({\bm{w}};{\bm{\xi}})]$. ∎
These limiting behaviors (Lemmas˜3.3 and 3.4) are revealing. As [NGS23a] does not characterize the worst-case distributions for (OR-WDRO) in a way that demonstrates reduced weighting of outliers in ${\mathbb{P}}_{\text{train}}$}, the convergence to CVaR-like measures suggests that OR-WDRO may inherently assign significant weight to data points that yield large losses. This is a characteristic of CVaR as previously discussed and contrasts with the goal of robust statistics to downweight or ignore outliers when learning from contaminated training data, especially if there is little prior knowledge about the covariance of the reference distribution ${\mathbb{P}}_{0}$ (thus it makes sense to set $\sigma\to\infty$). It is also clear that the limit behavior of (OR-WDRO) does not recover the robust statistics formulation of data contamination.
The intuition that OR-WDRO is not ideally suited for handling TV-contamination in the training set (a weaker model than the $\epsilon$-strong contamination considered in our work) can be further explored by setting $\rho=0$. In this case, the OR-WDRO problem simplifies to
$\min_{{\bm{w}}}\max_{{\mathbb{Q}}:\operatorname{TV}({\mathbb{Q}},{{\mathbb{P}}_{\text{train}}})\leq\epsilon}\mathbb{E}_{{\bm{\xi}}\sim{\mathbb{Q}}}[\ell({\bm{w}};{\bm{\xi}})].$
If the true clean distribution ${\mathbb{P}}_{0}$ is within this TV-ball, the objective robustifies against the worst case within this ball, which leads to the CVaR behavior. This approach, while providing a guarantee over the TV-ball, does not necessarily leverage more specific structural information about ${\mathbb{P}}_{0}$ beyond its containment. Although [NGS23a] use structural assumptions like moment constraints for their tractable variant, these are aids for computation rather than intrinsic parts of the original definition (OR-WDRO). Similar TV-robust formulations have also been noted for their sensitivity to training data outliers [BV22a, Remark 2.5]  [^7] Remark 2.5 in [BV22a] distinguishes their LP-DRO approach (which can amplify the influence of high-loss data points in training data) from traditional robust statistics, noting that the latter’s strategy of outlier identification and suppression often relies on structural assumptions about the true distribution ${\mathbb{P}}_{0}$. They position their work, which makes no such assumptions about ${\mathbb{P}}_{0}$, in contrast to [NGS23a], implying the latter operates under, or would require, such structural assumptions to effectively handle statistical outliers. We observe, however, that the fundamental OR-WDRO formulation (OR-WDRO) from [NGS23a] does not explicitly incorporate these structural conditions on ${\mathbb{P}}_{0}$ into its definition. Consequently, as our analysis indicates (e.g., Lemma 3.3), the OR-WDRO formulation (OR-WDRO) can also exhibit sensitivity to training data outliers, a characteristic effectively shared by the LP-DRO approach described in [BV22a]..
@@
${{\mathbb{P}}_{\text{train}}}={\mathbb{P}}_{0}$ $W_{p,\epsilon}({{\mathbb{P}}_{\text{test}}},{\mathbb{P}}_{0})\leq\rho$ OR-WDRO [NGS23a]
$W_{p,\epsilon}({{\mathbb{P}}_{\text{train}}},{\mathbb{P}}_{0})\leq\rho$ [^8] Training data in [PP24a] can simultaneously be $\epsilon$}-corrupted and have Wasserstein perturbation. ${{\mathbb{P}}_{\text{test}}}={\mathbb{P}}_{0}$ Local & global corruption [NGS24a, PP24a]
@@
Another approach that, like ours, simultaneously addresses both training data contamination (modeled via $\operatorname{TV}({\mathbb{P}}_{0},{{\mathbb{P}}_{\text{train}}})\leq{\epsilon}$) and distributional shifts at test time is the Distributionally Robust Outlier-aware Optimization (DORO) framework proposed by [Zha+21a]. DORO defines its objective, coined as DORO risk, as the minimum $f$-divergence-based DRO risk (with shift parameter $D_{f}({{\mathbb{P}}_{\text{test}}},{{\mathbb{P}}_{\text{train}}})\leq\rho$) among all possible $(1-\epsilon)$-subpopulations of the training data. While the authors provide theoretical guarantees for the DORO risk minimizer, their proposed algorithm lacks both error bounds and computational complexity guarantees; as such, it remains unclear whether the DORO risk minimizer can be approximated provably—let alone efficiently; see further discussion and a counterexample in Section˜B.1.
The work by [Wan+24a] introduces a DRO framework defined by a novel ambiguity measure inspired by Unbalanced Optimal Transport (UOT). They propose that UOT-based distance is inherently more resilient to outliers and their formulation subtracts a penalty term based on prior knowledge to further down-weight contaminated data. While [Wan+24a] used this approach to handle contaminated training data and provided empirical results demonstrating improved robustness in such settings compared to standard Wasserstein DRO and OR-WDRO, they did not establish explicit theoretical guarantees on estimation error. More importantly, from the modeling perspective of (2), its mechanism of defining an ambiguity suggests that the outliers it robustifies against are perhaps best characterized in the testing distribution (${\mathbb{P}}_{\text{test}}$}), as discussed previously for [NGS23a] and in Section˜3.1.
In the sequel, we use data contamination or corruptions to specifically indicate the existence of outliers in the training data, whereas outliers might exist in either training or testing data.